write the function for a line y-intercept of 19 and goes through the point (17,-83)
Found 3 solutions by richwmiller, Cromlix, stanbon:
Answer by richwmiller(17219)

you have two points
(0,19) and (17,-83)
if you find the slope you can fill in the equation
y=mx+b
where b is the y intercept
so we have
y=mx+19
where m is the slope
-83-19 over 17-0
-102/17
y=(-102/17)x+19

Answer by Cromlix(4381)

Using the coordinates (0,19) and (17,-83)
Equation for slope (m) = y2 - y1/x2 - x1
= -83 - 19/17 - 0 = -102/17
Equation of a line:
y = mx + c
y = -102/17x + 19.
Hope this helps.
:-)

Answer by stanbon(75887)

write the function for a line y-intercept of 19
and goes through the point (17,-83)
----------
Form:: y = mx + b
-83 = m(17) + 19
-----
-102 = 17m
m = -102/17
m = -6
=======
Equation:
y = -6x + 19
===============
Cheers,
Stan H.
===============